A car engine whose output power is 205 hp h p operates at about 18 % % efficiency. Assume the engine's water temperature of 94 ∘C ∘ C is its cold-temperature (exhaust) reservoir and 495 ∘C ∘ C is its thermal "intake" temperature (the temperature of the exploding gas-air mixture). How much heat, in joules and in kcal, is exhausted to the air in 1.0 hh.
Answer #1

still any doubts feel free to ask at least before giving downvote hope you understand my concerns

Please like
Answer #2

SOLUTION 


Output power  = 205*746 = 152930 W 


Energy output in 1 hr (3600 s) = 152930 * 3600 W-sec = 550548000 J. (1 w-sec = 1 J)


Energy input = Output / 0.18 = 550548000 / 0.18 = 3058600000 J


Energy lost = heat exhausted to the air = 3058600000 -  550548000 = 2508052000 J

= 2.508*10^9 J (ANSWER)


Heat exhausted to air in kCal = 2.508*10^9 *  2.39*10^(-4)  (i J = 2.39*10^(-4) kCal)

= 5.99424*10^5 kCal (ANSWER).

answered by: Tulsiram Garg